OpenStudy (jamesj):

Here's an interesting integral: the primitive of 1/sqrt( x(1-x) )

OpenStudy (jamesj):

\[ \int \frac{dx}{\sqrt{x(1-x)}} \] I have a solution for it, but would interested to see how others approach the problem. This is one case (among others!) where WolframAlpha is actually not very helpful.

OpenStudy (anonymous):

I am not sure if what I have in mind is somehow useful for this kind of problem, but I was considering expending the term under the radical and then making a perfect square out of it. Afterwards a trig substitution.

OpenStudy (anonymous):

Yep, completing the square would be the most apparent route.

OpenStudy (jamesj):

ok ... so work it through

OpenStudy (anonymous):

I am not too sure about it yet. Some of the signs bother me still

OpenStudy (jamesj):

Yes, the signs are bizarre and things become problematic. I had to take things into complex numbers to make it easy to work through at first and then do some slightly non-standard (for an integral problem) manipulation to get back to a real function.

OpenStudy (anonymous):

\[\int \frac{dx}{\sqrt{x^2-x}}\] x² - x x² - x + 1/4 - 1/4 (x - 1/2)² - 1/4 \[\int \frac{dx}{\sqrt{(x-\frac{1}{2})^2-\frac{1}{4}}}\] Let x - 1/2 = 1/2 secu dx = 1/2 secu tanu du \[\int \dfrac{\frac{1}{2}\sec u\tan u}{\sqrt{(\frac{1}{2}\sec u)^2 - \frac{1}{4}}}du\\ \frac{1}{2}\int \dfrac{\sec u\tan u}{\sqrt{\frac{1}{4}\sec^2 u - \frac{1}{4}}}du\\ \int \dfrac{\sec u\tan u}{\sqrt{\sec^2 u - 1}}du\\ \int \dfrac{\sec u\tan u}{\sqrt{\tan^2 u}}du\\ \int \dfrac{\sec u\tan u}{\tan u}du\\ \int \sec u\;du\] And from there, it's all a matter of subbing back.

OpenStudy (anonymous):

For the Radical: \[\Large \sqrt{-x^2+x-\frac{1}{4}+\frac{1}{4}} = \sqrt{-(x-\frac{1}{2})^2-\frac{1}{4}} \]

OpenStudy (anonymous):

So I see where the signs get complicated.

OpenStudy (anonymous):

Oops, looks like I made a mistake with the sign... Just a sec.

OpenStudy (jamesj):

My Sith friend, it is \( \sqrt{x - x^2} \), which does make a difference.

OpenStudy (anonymous):

in fact, it makes all the difference (-: ! I would go with a complex substitution now too. Let me see.

OpenStudy (anonymous):

Here's the change to my original reply: In the denominator, under the radical, you'd have -x² + x + 1/4 - 1/4 -(x² - x + 1/4 - 1/4) -((x - 1/2)² - 1/4) 1/2 - (x - 1/2)² Let u = x - 1/2 du = dx So the integral becomes \[\int \dfrac{du}{\sqrt{\frac{1}{4}-u^2}}\] Then, let u = 1/2 sint du = 1/2 cost dt I think this should fix that.

OpenStudy (anonymous):

Unless I made some minor algebraic mistake again...

OpenStudy (anonymous):

Looks like I did, but nothing really changes. In the fourth line, the 1/2 should be 1/4.

OpenStudy (jamesj):

so your final answer comes out as (1/2).arcsin(2x-1), yes?

OpenStudy (jamesj):

Nice.

OpenStudy (anonymous):

I haven't worked it out completely, but that's what it looks like it'll turn out to be.

OpenStudy (jamesj):

Yes, that's right. From there you can deduce this interesting result as well \[ \int_0^1 \frac{dx}{\sqrt(x(1-x)} = \pi \] *** More methods welcome.

OpenStudy (anonymous):

I concur, Just that my substitution now (on a second tryI) was different. \[\Large \alpha = \sin^{-1}(2x-1) \]

OpenStudy (jamesj):

\[ \int_0^1 \frac{dx}{\sqrt{x(1-x)}} = \pi \]

OpenStudy (jamesj):

right, the (1/2) I had there was wrong.

OpenStudy (anonymous):

Well, your title wasn't promising too much @JamesJ. A very interesting integral.

OpenStudy (jamesj):

Here's my solution for the record. There's a slight bug towards the end with a minus sign. The final answer should be \[ I = -2 \arctan\sqrt{\frac{1}{x} - 1} \] http://openstudy.com/study#/updates/5112b975e4b0e554778a94c3

OpenStudy (jamesj):

Which also suggests one of those truly arcane identities: \[ \arcsin(2x - 1) + 2\arctan\sqrt{\frac{1}{x} - 1} = \frac{\pi}{2} \]

OpenStudy (rogue):

I evaluated the integral via substitution. It works great. :D\[\int\limits\limits \frac {dx}{\sqrt{x} \sqrt {1 - x}}, u = \sqrt {1-x}, -2du = \frac {dx}{\sqrt{1-x}}, \sqrt {x} = \sqrt {1 - u^2}\]\[\int\limits \frac {-2du}{\sqrt{1 - u^2}} = -2 \sin^{-1} (u) = -2 \sin^{-1} (\sqrt{1-x})\]

OpenStudy (jamesj):

Yes, that's neat too. Good stuff.